### v1

# Evidence-Decision-Feedback: Theory-Driven Adaptive Scaffolding for LLM Agents

Abstract
Multi-agent LLM architectures offer opportunities for pedagogical agents to help students construct domain knowledge and develop critical-thinking skills, yet many operate on a “one-size-fits-all” basis, limiting their ability to provide personalized support. To address this, we introduce Evidence-Decision-Feedback (EDF), a theoretical framework for adaptive scaffolding using LLMs. EDF integrates elements of intelligent tutoring systems and agentic behavior by organizing interactions around evidentiary inference, pedagogical decision-making, and adaptive feedback. We instantiate EDF through Copa, an agentic collaborative peer agent for STEM+C problem-solving. In an authentic high school classroom study, we show that EDF-aligned interactions align feedback with students’ demonstrated understanding and task mastery; promote gradual scaffold fading; and support interpretable, evidence-grounded explanations without fostering overreliance.
1 Introduction
Large language model (LLM) agents afford dialogic support that aligns with learning theories emphasizing knowledge construction through social interaction, including Social Cognitive Theory [1] (SCT), social constructivist perspectives, and the Zone of Proximal Development (ZPD) [30]. However, recent studies have examined the extent to which LLMs support (or hinder) learning, with emerging findings raising concerns: Kosmyna et al. [18] reported 83% of students receiving ChatGPT assistance during essay writing were unable to quote their own work, and Zhou et al. [32] found GPT-4 feedback on compiler errors improved performance and persistence during assisted tasks but not after feedback was withdrawn. These agents can encourage overreliance, thereby impeding sustained learning and critical thinking. As such, Borchers et al. [3] concluded that LLM tutoring in its current form is “unlikely to produce learning benefits rivaling known-to-be-effective ITS tutoring.”
Prior research has also shown that students often attempt to “game” intelligent learning systems by prioritizing task completion and performance metrics over deeper conceptual understanding [5]. Such behaviors can create the appearance of mastery without corresponding gains in underlying conceptual knowledge, diverging from learning-theoretic aims that emphasize sense-making, reflection, and principled scaffolding. These concerns highlight the need for pedagogical agents that do more than optimize task performance, motivating a growing call within the AIED community for LLM-based agents to support and encourage critical thinking while remaining grounded in learning theory [27, 8].
Such systems must be personalized to students’ diverse needs while generating feedback interpretable to stakeholders, who may be wary of adopting systems whose reasoning is opaque or inadequately explained [11]. The AI community has recently embraced agentic AI, where agents autonomously plan actions, reason over data, and execute tasks to achieve defined goals, often within multi-agent architectures [28]. These systems are well-suited to educational contexts, as agent reasoning can support adaptation and interpretability by grounding decisions in learner-specific data and linking feedback to established learning theories. However, despite the proliferation of LLM agents in AIED research and practice, there remains a lack of LLM-era frameworks that operationalize agentic systems in systematic, pedagogically meaningful ways.
There is a critical need for LLM-era frameworks that are (1) adaptive to diverse student needs, (2) effective in supporting student understanding, and (3) interpretable to stakeholders. In response, we introduce a theoretical framework for adaptive scaffolding—Evidence-Decision-Feedback (EDF; Section 3)—that draws on foundational theories including Evidence-Centered Design (ECD) [21], Stealth Assessment [26], SCT, ZPD, and social constructivist perspectives to guide the design of pedagogically grounded LLM-based agents that support student learning in real time.
We developed an agentic, collaborative peer agent, Copa, to instantiate EDF by mapping each EDF component to concrete elements of Copa’s multi-agent architecture (Section 4). We deploy Copa in an authentic high school science classroom with $n=33$ dyads, where students learn physics and computing by building computational models, with the agent supporting inquiry, exploration, and critical thinking during problem-solving. Our findings suggest that EDF supports personalized, effective, and interpretable adaptive scaffolding without fostering overreliance on the agent, advancing the development of theory-aligned, agentic LLM systems in AIED.
2 Background
Before LLMs, intelligent tutoring system (ITS) frameworks like Adaptive Control of Thought–Rational (ACT-R) and Knowledge-Learning-Instruction (KLI) focused on modeling human cognition to predict task performance and align instructional design with cognitive processes. While this enabled precise model-tracing feedback, these frameworks were not well-suited to open-ended learning tasks and struggled to address the complexity and opacity of LLM systems.
Recent work has sought to improve the interpretability of LLM-based feedback by incorporating explanatory reasoning into agent responses, often using chain-of-thought (CoT), to align outputs with prompt instructions [11, 6]. While this enables LLMs to reconcile input data with instructional constraints before generating feedback, such interpretability is typically limited to surface-level explanations, leaving the underlying decision-making processes opaque.
With agentic and multi-agent architectures, decision-making can be modularized, and interactions among tools and agents can be traced back to underlying decision-making processes. Li et al.’s EduMAS framework [19] employs Concept Clarifier and Experimentation Mentor agents that collaborate to support comprehensive learning experiences. Chu et al.’s GradeOpt system [4] integrates Grader, Reflector, and Refiner agents to improve alignment with human experts. Dai et al.’s Agent4EDU framework [12] uses agentic workflows to enable targeted instructional support. However, these systems provide limited guidance on grounding agent scaffolding in pedagogy.
Many existing agents also offer limited adaptive dialogic support. Google’s LearnLM-Tutor provides scaffolding informed by stakeholder co-design [17] but does not personalize its interactions. Thomas et al. [29] ground system scaffolding in feedback theory but without dialogic interaction. Other studies similarly employ LLMs to deliver single-shot feedback, lacking the interactive and adaptive qualities central to effective dialogic learning [23, 31, 20].
The current LLM agent landscape in AIED reveals two persistent gaps: (1) a lack of principled mechanisms for personalizing scaffolding to individual learners and (2) insufficient support for interpreting how agent decisions are derived from learner evidence. While research has explored adaptive scaffolding prior to LLMs [2, 22] and in task-specific LLM applications [8], no comprehensive framework exists that addresses the following overarching question: How can we support LLM-based adaptive scaffolding that is personalized, effective, and interpretable?
3 Theoretical Framework
Figure 1: Evidence-Decision-Feedback (EDF) framework.
Over four years of participatory design (PD) with students, teachers, and learning scientists, we synthesized surveys, interviews, focus groups, pilot studies, and collaborative design activities into a framework aligned with established theory [10, 5, 9, 8, 11, 13]. We present Evidence-Decision-Feedback (EDF), a principled approach to adaptive scaffolding for LLM-based pedagogical agents that addresses the limitations of existing frameworks discussed in Section 2.
EDF comprises three semi-autonomous modules that provide a blueprint for operationalizing adaptive support (shown in Figure 1). Using ECD and Stealth Assessment, the Evidence module enables stakeholders to specify the student data to be monitored to continuously update a learner model. This provides mastery evidence and is used by the Decision module to determine pedagogical intent (e.g., boosting self-efficacy) by generating a dialogue policy that is consistent with SCT and within the ZPD. The Feedback module operationalizes this policy into adaptive scaffolding, ensuring it reflects social constructivist principles. When a student responds, the response is stored as dialogue evidence and used to update the learner model before the next conversation turn.
Gray dotted arrows in Figure 1 denote interpretability, a core tenet of EDF. We define interpretability as stakeholders’ ability to understand why agent output was generated by examining how observable input data is reflected in evidentiary representations and policy explanations. Using CoT prompting, we elicit the LLM’s reasoning within each module so that agent feedback can be traced to dialogue policy selection, policy-to-evidentiary inferences, and evidence-to-input data. This provides a transparent account of how student input data is transformed into agent output, offering insight into agent decision-making.
4 Copa
To operationalize EDF, we developed Copa—a Collaborative Peer Agent for STEM+C learning. Copa is designed as a knowledgeable peer—a persona selected by students who prefer the emotional support of a peer and domain competence of an expert [5]. This role encourages inquiry, exploration, and critical thinking while minimizing the dependency risk and self-efficacy loss often associated with expert tutors. Copa is agentic, using sub-agents to gather evidence, reason about learners’ needs, and provide scaffolding within the EDF framework.
Copa reasons over student log and chat data, selects actions, and generates feedback while under constraints. This aligns with Jurenka et al.’s [17] “see what the student sees” agent design principle and prior work suggesting bounded autonomy [24]—which balances agent initiative with pedagogical guardrails, e.g., human-in-the-loop methods [13]—represents an ideal design point to support learning while preserving trust and interpretability [11, 8]. Rather than relying on unconstrained long-horizon planning, Copa’s behavior is structured around the three EDF modules. Copa’s four-agent architecture is grounded in PD, which identifies appropriate agent behavior and the data to be considered, and in prior work demonstrating that modular, multi-agent systems improve transparency and task decomposition in educational settings [19, 4, 12].
Figure 2: Copa, pictured within the C2STEM learning environment.
Copa is embedded within the computer-based C2STEM [15] learning environment (see Figure 2). C2STEM targets 1- and 2-D kinematics: students work in dyads on various tasks, building computational models to simulate the motion of objects such as a truck (1-D) and a drone (2-D) using kinematic equations and block-based code, and interacting with Copa at their discretion. Each task is open-ended: there is no single solution, and each may be solved in myriad ways.
C2STEM interactions generate rich log data to fuel the agent’s reasoning, including logged actions (raw action sequences like adding/editing/removing blocks), model state (the code blocks on screen), and task context (the model component the student is currently addressing, e.g., variable initialization). Logged actions are mapped to processed actions to create a semantically meaningful representation for the LLM (e.g., set_block1234_vel_4 $\rightarrow$ Set Velocity to 4 m/s).
Log data also enables us to track students’ progress by defining rubric criteria that must be met for each task, even though tasks are open-ended. By translating students’ code into a canonical form, we can use the model state to determine if students have properly initialized variables, for example. Importantly, we distinguish between students’ understanding—their verbalized knowledge to the agent demonstrating what they know—and their task mastery—performance measured as the percentage score achieved for a given C2STEM task.
Figure 3: Copa and its four sub-agents. Colors correspond to the EDF modules presented in Figure 1. Gray boxes = data stores. Thick edges = two-way relationships.
Shown in Figure 3, Copa comprises four specialized sub-agents operating over a shared learner model that is continuously updated through activity logs and agent conversations. This architecture directly maps onto the EDF modules via its color-coded sub-agents, with particular emphasis on the Evidence module (blue), which comprises three asynchronous agents that transform noisy log data into a coherent learner model. All four agents additionally emit evidence with their outputs, using CoT to justify decisions and ensure interpretability.
The StrategyAgent analyzes temporal action patterns to infer cognitive problem-solving behaviors, distinguishing productive strategies (e.g., TINKERING: systematic trial-and-error) from haphazard approaches (e.g., DEPTH-FIRST ENACTING: randomly building without periodic testing). The AssessmentAgent triangulates the current model state and strategy with past learner states (e.g., DEBUGGING or STRUGGLING) to generate an updated learner state. The KnowledgeAgent compares the model state to an expert reference to identify conceptual gaps within the ZPD and retrieves aligned domain knowledge (i.e., kinematics/computing information) via retrieval-augmented generation (RAG).
The DialogueAgent is the central orchestrator, integrating the Decision (green) and Feedback (purple) modules. These modules must operate synchronously, so housing them within a single agent minimizes latency, enabling asynchronous reasoning via the Evidence agents without interrupting conversational flow. Upon receiving a student query, the DialogueAgent identifies the students’ dialogue state (e.g., DEMONSTRATES_UNDERSTANDING, where students verbalize conceptual knowledge to Copa). The agent then synthesizes the dialogue state and learner model to determine pedagogical intent and generate a dialogue policy, e.g., PROBE_UNDERSTANDING (i.e., students want a direct solution or demonstrate a misconception, prompting Copa to first probe their knowledge), SUGGEST_ACTION (i.e., take a concrete action in C2STEM), or PUSH_LIMIT (i.e., push students to extend and deepen their understanding). This policy is then used by the DialogueAgent to generate a talk move, i.e., a response to students.
We evaluated LLMs from the Gemini, Claude, and GPT families and selected GPT for optimal performance. Asynchronous Evidence agents use the high-reasoning capabilities of GPT-5, while the synchronous DialogueAgent employs GPT-5-Chat to support low-latency interaction. Prompts were developed using a human-in-the-loop procedure from prior work, CoTAL [11], which combines CoT prompting with iterative refinement (active learning) to link student input data to teacher and researcher instructions. Prompts were refined across four hour-long sessions with 10-15 researchers, who interacted with Copa using C2STEM and provided feedback that informed subsequent revisions.
All agents explicitly operationalize learning theory. For example, the KnowledgeAgent estimates students’ ZPD and identifies the “immediate next step” needed based on task progress, while the DialogueAgent is instructed to “promote self-efficacy through encouragement and praise” (SCT) and ask “probing questions” (Social Constructivism). All task rubrics, as well as agent prompts, features, and their values, are detailed in our Supplementary Materials [^1] https://github.com/claytoncohn/AIED26_Supplementary_Materials.
5 Evaluation
To evaluate Copa, we investigate the following Research Questions:
  1. 1.
    Do Copa’s scaffolds appropriately adapt as students increase task mastery?
  2. 2.
    Is students’ demonstrated understanding of the material aligned with their task mastery when interacting with Copa?
  3. 3.
    Do students rely less on Copa as they achieve greater task mastery?
  4. 4.
    Is Copa’s feedback interpretable with respect to its evidentiary inferences?
In collaboration with five STEM educators and cognitive scientists (hereafter, “the educators”), we deployed Copa in an authentic high school science classroom in the southeastern United States with $n=33$ sophomore dyads [^2] We conducted all analyses at the dyad level (one student-pair per computer). (ages 15-16; 48% male/female, 2% no response; 68% White, 28% Hispanic, 20% Black, 12% Asian, 4% Native American) [^3] Several students identified as multiple ethnicities. participating in the C2STEM kinematics curriculum over six weeks. Sessions met once weekly for two hours and were taught by this paper’s lead author with oversight from the educators.
Students were first introduced to Copa five months prior to the study through a “teaser” activity that covered its pedagogical role, design rationale, and how it differs from tools such as ChatGPT; rather than providing direct solutions (ChatGPT), Copa supports students in constructing their own understanding. During the study, students completed two warm-up tasks to gain familiarity with Copa. The analyses in this paper focus on three subsequent tasks following this acclimation period. Each session began with instruction on 1- and 2-D kinematics, linking physics concepts (e.g., position, velocity, acceleration, $\Delta t$) with computational constructs (e.g., variable initialization and updates, loops, and conditionals), followed by computational modeling activities in C2STEM.
For this study, we collected 7,017 logged environment actions (mean $\approx 213$ per dyad per session), and 238 student-agent conversation turns (mean $\approx 7$ per dyad per session) across the three sessions, along with the CoT reasoning (i.e., “evidence” in Figure 3) associated with each agent decision. Together, these data streams constitute the primary dataset used for analysis in this paper, with all inferences manually triangulated using students’ conversations and screen recordings. Students completed an anonymous semi-structured exit survey. All findings were presented to three of the participating educators during a semi-structured Zoom discussion, providing additional feedback. Study participants provided informed consent/assent, student data were anonymized prior to analysis, and all procedures were approved by Vanderbilt University’s IRB.
5.1 RQ1: Scaffolding Adaptivity
As students achieve greater task mastery, Copa’s scaffolding is expected to adapt by gradually fading in accordance with ZPD principles. At the outset, Copa should prioritize probing students’ understanding to assess their current knowledge, approximate their ZPD, address prior misconceptions, and encourage critical thinking. As students deepen their understanding, Copa should look ahead, offering concrete suggestions to help them extend their reasoning beyond their current level of competence.
To evaluate Copa’s adaptivity (RQ1), we examined how its dialogue policies evolved within each session by analyzing the frequency with which policies were selected relative to quintiles of students’ task mastery percentages (see Table 1). We used quintiles for more stable comparisons of agent behavior across meaningful phases of learning. Policy frequencies were normalized as a percentage of the total dialogue within each dyad’s session, and then averaged across mastery quintiles. For each dialogue policy, we computed Spearman’s $\rho$ (due to the ordinal nature of the mastery data) to assess the strength of the relationship between shifts in dialogue policy and task mastery.
Dialogue Policy Spearman’s $\rho$ Trend $p$-value
PROBE_UNDERSTANDING 0.34-0.34 Decreasing $0.034$
SUGGEST_ACTION $0.33$ Increasing $0.039$
PUSH_LIMIT $0.42$ Increasing $0.007$
Table 1: Dialogue policy adaptation across task mastery quintiles.
As shown in Table 1, Copa significantly [^4] Throughout the paper, statistical significance is assessed at $\alpha=0.05$. reduced its use of the PROBE_UNDERSTANDING policy as mastery increases ($\rho=-0.34$, $p=0.034$), shifting toward implementation- and expansion-oriented support: SUGGEST_ACTION ($\rho=0.33$, $p=0.039$) and PUSH_LIMIT ($\rho=0.42$, $p=0.007$) increase in relative frequency. This demonstrates a pedagogical transition in which Copa moves from eliciting and assessing conceptual understanding to supporting the application of that understanding through model construction as mastery increases. This reflects ZPD alignment, as Copa shifts assistance toward challenges that extend learners beyond their independently-demonstrated competence, indicating Copa’s scaffolding appropriately adapted as task mastery increased.
5.2 RQ2: Understanding-Mastery Alignment
Students’ prioritization of task completion over conceptual understanding can present an overly optimistic picture of their learning, as students may appear to progress without being able to explain their code. From a social constructivist perspective, learning is evidenced not only by task performance but by learners’ ability to articulate and negotiate meaning through dialogue and explanation. Evaluating whether students’ task mastery is meaningfully connected to their verbalized understanding while interacting with Copa is, therefore, paramount.
To address RQ2, we examined the relationship between students’ response success rates—defined as the proportion of instances students correctly explained their code via a DEMONSTRATES_UNDERSTANDING dialogue state immediately following a PROBE_UNDERSTANDING dialogue policy—and students’ task mastery deciles. Deciles provide a finer-grained analysis of how students’ verbally demonstrated understanding varies across incremental levels of task mastery in C2STEM to assess alignment between the two. Data were normalized by averaging response success rate percentages for each dyad to mitigate interaction frequency bias. We again used Spearman’s $\rho$ due to the ordinal nature of the mastery deciles.
Figure 4: (a) Correlation between probing success rate and mastery deciles (RQ2).
(b) Correlation between student requests for agent support and mastery deciles (RQ3).
Figure 4a shows the success rate [^5] Success rate for mastery >0.7>0.7 was not calculated due to a lack of PROBE_UNDER-
STANDING policies: Copa chose to not probe knowledge once mastery was sufficient, as students stopped asking for direct answers and demonstrating misconceptions.
increases significantly as mastery increases ($\rho=0.40$, $p=0.014$), suggesting students are not merely “gaming the system”: mastery gains are accompanied by an improved ability to verbally reason about their code and respond correctly to Copa, i.e., C2STEM progress is meaningfully aligned with conceptual understanding rather than superficial task optimization.
Students’ environment actions advanced their model rather than deteriorated it, at approximately a 3:1 ratio, following agent interactions. The DEMONSTRATES_UNDERSTANDING dialogue state also showed a significant percentage increase as mastery improved ($\rho=0.34$, $p=0.030$), evidencing students’ increasing ability to articulate their reasoning as their problem-solving progressed. From a social constructivist perspective, articulating reasoning through dialogue reflects learners’ active construction of knowledge via interaction. Together, these findings demonstrate that students’ verbalization of their conceptual understanding aligned with their C2STEM task mastery.
5.3 RQ3: Student Reliance on Agent
Students’ control over how and when they interact with Copa risks overreliance. Ideally, as students approach task mastery, they would engage with Copa less frequently, reflecting increasing independence and reduced reliance on external support. Pursuant to SCT, this transition reflects the development of self-regulation and self-efficacy skills. However, prior work suggests this is not always the case, and students may develop a dependence on agent feedback that ultimately impedes their learning. This necessitates evaluating whether students use Copa in a manner consistent with productive support or continue to rely on the agent despite having developed sufficient task knowledge.
To evaluate the appropriateness of students’ reliance on Copa (RQ3), we examined how students’ voluntary engagement with the agent changed as they progressed through task mastery deciles. We computed the percentage of agent support (i.e., interactions) across mastery deciles to quantify the proportion of Copa turns per session. We again used deciles to provide fine-grained resolution over students’ progression toward mastery. To control for bias introduced by imbalanced interaction volumes, we normalized the data by averaging the agent-support proportion for each dyad across deciles, ensuring that high-frequency dyads did not exert disproportionate influence. We again used Spearman’s $\rho$ due to the ordinal nature of the mastery deciles.
Figure 4b shows students’ requests for support are highest early on, with approximately 59% of support occurring while mastery is low (<40<40%). As mastery increases, we observed a significant drop in requests for agent feedback (Spearman’s $\rho=-0.26$, $p<0.001$): students increasingly relied on their own knowledge, critical thinking, and collaboration with their partner rather than on agent support as mastery improved, consistent with developing learner autonomy, self-regulation, and self-efficacy—central SCT tenets.
5.4 RQ4: Feedback Interpretability
To trust agents, teachers, students, parents, and administrators must understand how they make decisions; otherwise, they risk limited adoption. EDF’s emphasis on interpretability aligns with ECD and Stealth Assessment, which require inferences about student knowledge be grounded in observable evidence collected unobtrusively during interaction. Agent feedback must be traceable to the dialogue policy that generated it, the evidence identified by the agent, and ultimately to the raw student data. To assess the interpretability of Copa’s feedback (RQ4), we decompose Copa’s internal reasoning chain (the agents’ “evidence” outputs in Figure 3) into the three links between EDF modules:
$\text{Student Input Data}\xrightarrow{\text{Link 1}}\text{Evidence}\xrightarrow{\text{Link 2}}\text{Decision}\xrightarrow{\text{Link 3}}\text{Feedback}$
Link 1, referred to as Grounding, measures the extent to which Copa’s Evidence extraction reflects student input data. We evaluate Grounding using keyword recall, defined as the percentage of semantically meaningful tokens from student log data that appear in the extracted evidence. We apply Porter stemming to account for morphological variation (e.g., accelerates $\approx$ acceleration), retaining only tokens longer than three characters with at least one alphabetic character. Significance is assessed via a permutation test ($n=100$), comparing observed log-evidence pairs against randomly shuffled baselines.
Link 2, referred to as Alignment, assesses semantic coherence between Copa’s Decision (dialogue policy selection) and extracted Evidence. We compute sentence-level semantic similarity using Sentence-BERT embeddings (SBERT; all-MiniLM-L6-v2)—selected for its balance of efficiency and performance—by encoding Copa’s dialogue policy and dialogue state summaries, then comparing them via cosine similarity. Significance is evaluated via a permutation test (n=1000n=1000) [^6] Larger $n$ due to fewer utterances relative to environment actions in Link 1., comparing mean similarity of true evidence-policy pairs to shuffled baselines. Link 3, referred to as Faithfulness, evaluates whether Copa’s final Feedback adheres to its Decision. We assess faithfulness by again computing the SBERT-based cosine similarity between the dialogue policy and talk move, using the same permutation testing procedure as Link 2.
Trace Component Metric Actual Random p-value
Grounding (Data $\rightarrow$ Evidence) Keyword Recall 0.43 0.21 $p<0.001$
Alignment (Evidence $\rightarrow$ Decision) SBERT Similarity 0.64 0.39 $p<0.001$
Faithfulness (Decision $\rightarrow$ Feedback) SBERT Similarity 0.48 0.24 $p<0.001$
Table 2: Interpretability results for all three links.
Table 2 shows that all three links exhibit statistically significant non-random structure ($p<0.001$), with similarity scores for true pairs approximately doubling those of the baselines. The Grounding score indicates Copa’s evidence extraction captured approximately twice as many meaningful tokens from the logs relative to the baseline (0.43 vs. 0.21). The Alignment score of 0.64 indicates strong semantic coherence between the extracted evidence and the dialogue policy selections, well above the baseline of 0.39. The Faithfulness score of 0.48 demonstrates that Copa’s generated responses remain semantically aligned with the dialogue policy, also outperforming the baseline (0.24). Viewed through the lens of ECD and Stealth Assessment, these results indicate Copa’s decisions are interpretable and grounded in explicit evidentiary reasoning.
5.5 Student and Teacher Perceptions
Post-study student survey responses revealed students responded positively to Copa overall, agreeing that “the questions Copa asked were appropriate” (mean =3.81=3.81 on a five-point Likert scale) and Copa “asks questions that made me think about my model” (mean =3.88=3.88). This suggests Copa successfully implemented its intended design goal of prioritizing critical thinking over direct solutions.
This behavior was not always appreciated. Ratings for perceived agent understanding (“I feel like Copa understood what I was doing”) and feedback utility (“I feel like Copa’s suggestions were useful”) were lower (mean =2.69=2.69 and $2.77$, respectively), indicating a disconnect between Copa’s pedagogical intent and students’ desire for answers. One group reported that Copa was “broken” and expected direct feedback rather than ZPD-aligned hints. This illustrates a broader tension between principled, scaffolded support and students’ expectations.
Findings aligned with educators’ classroom experiences. They noted that (1) students become frustrated when they do not receive immediate answers (“I agree with your conclusion…students wanting immediate answers”); (2) this expectation is shaped by their prior experiences with ChatGPT (“ChatGPT [is] so desperate to provide help…they’ve already become accustomed”); and (3) a persistent tension exists between instructional goals and student wishes, as students often “just [want]…the right answer…check the box and move on.”
One educator observed that student frustration was most pronounced due to repetitive interaction loops in which Copa withheld progress until understanding was demonstrated—particularly when students declined to act on feedback—noting that, while pedagogically intuitive, such questioning may not always be optimal [25]. Another viewed this positively, noting this behavior “gives Copa more of a human response”. Educators emphasized that students hold different expectations for agents than humans, underscoring the importance of student buy-in. To support this, one educator suggested sharing prior study findings with new cohorts at the outset to communicate Copa’s benefits and having Copa explain why students are stuck in such loops during interaction as a strategy for improving transparency and engagement.
6 Discussion and Conclusions
This work introduces an LLM-era, pedagogy-explicit scaffolding framework—Evidence-Decision-Feedback (EDF)—that treats adaptation and interpretability as first-class design goals. Operationalized through Copa, we demonstrated that an EDF-driven agent can adapt feedback to students’ evolving mastery in ways that are interpretable, avoid overreliance, and support the verbalization of conceptual understanding, consistent with social cognitive and social constructivist principles. Our modular, agentic design tightly couples learning theory with system architecture and is evaluated in an authentic classroom setting, rather than solely through benchmarks or single-shot feedback.
Theory emphasizes that effective support should help students build understanding rather than just providing answers. However, consistent with prior work [7], our findings indicate that students do not always value this approach. Research also emphasizes the importance of learner agency for engagement [5], raising questions about the appropriate degree of autonomy when it conflicts with established pedagogical principles. Authors et al. [7] argue that student agency should be supported only to the point that it does not undermine critical thinking opportunities, a view supported by our findings. Overall, our results indicate that while EDF allows Copa to apply learning theories effectively, they reveal a tension between principled guidance and students’ preferences.
From a methodological perspective, our results emphasize adaptation as a first-class system capability, requiring continuous scaffold adjustments based on learner evidence rather than executing fixed strategies. Recent work on agentic AI failures found that agents often appear effective during testing but degrade in real use due to a lack of adaptation [16]. This reinforces the need to evaluate agentic systems in authentic classroom settings, where learners’ evolving behaviors, needs, and goals reveal limitations not visible in controlled benchmarks.
Practically, effective deployment of agentic systems in situ requires attention beyond system design. As Thomas et al. [29] note, the value of agent feedback ultimately depends on students’ willingness to engage with it. Our findings reinforce the need for the AIED community to invest in fostering student buy-in and trust, as meaningful learning benefits are unlikely if students disengage from or resist agent support. Deeper engagement is, therefore, a prerequisite for realizing the pedagogical potential of LLM agents in practice.
Our limitations include evaluating scaffold adaptation as a function of task mastery rather than a broader range of attributes, and the absence of an RCT due to institutional constraints. Reported relationships are correlational, not causal, and may be influenced by students’ prior knowledge and interaction dynamics. However, correlational evidence has been shown to be valuable for informing student performance, understanding, and the role of pedagogical agents in authentic settings [14]. We note (for reference only) that students using Copa exhibited higher normalized learning gains (0.53) than the prior cohort not using an agent (0.45) from the same school, curriculum, and demographics.
Overall, our work demonstrates that (1) agentic, multi-agent LLM systems can be designed to operate in principled, interpretable, and pedagogically meaningful ways; and (2) EDF offers a viable framework for operationalizing learning theory within LLM-based agents towards these ends. In future work, we will expand our adaptive scaffolding evaluation beyond task mastery to include factors like prior knowledge, self-regulation, self-efficacy, and social dynamics. We hope this work inspires further AIED research that moves beyond surface-level LLM interactions toward theoretically grounded, adaptive, and trustworthy educational agents that can meaningfully support learning at scale.
{credits}
6.0.1 Acknowledgements
The research reported here was supported by the Institute of Education Sciences (IES), U.S. Department of Education, through Grant R305C240010; and National Science Foundation (NSF) awards IIS-2327708 and DRL-2112635 as subawards to Vanderbilt University. The opinions expressed are those of the authors and do not represent the views of the Institute of Education Sciences, U.S. Department of Education, or National Science Foundation.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.